### v1

# Quasiprobability Thermodynamic Uncertainty Relation

Abstract
We derive a quantum extension of the thermodynamic uncertainty relation where dynamical fluctuations are quantified by the Terletsky–Margenau–Hill quasiprobability, a quantum generalization of the classical joint probability. The obtained inequality plays a complementary role to existing quantum thermodynamic uncertainty relations, focusing on observables’ change rather than exchange of charges through jumps and respecting initial coherence. Quasiprobabilities show anomalous behaviors that are forbidden in classical systems, such as negativity; we reveal that such behaviors are necessary to reduce dissipation beyond classical limitations and show that they are stronger requirements than that the state has quantum coherence. To illustrate these statements, we employ a model that can exhibit a dissipationless heat current, which would be prohibited in classical systems; we construct a state that has much coherence but does not lead to a dissipationless current due to the absence of anomalous behaviors in quasiprobabilities.
Introduction.— Finding universal lower bounds on the entropy production, which quantifies the irreversibility in the process, is one of the main tasks in nonequilibrium thermodynamics because they provide fundamental limitations beyond the second law of thermodynamics [1]. The thermodynamic uncertainty relation (TUR) is arguably the most crucial example, having been intensively studied over the last decade [2, 3]. It consists of dissipation, quantified by the entropy production rate (EPR) $\dot{\Sigma}$, and a general current’s strength $J_{X}$ and fluctuations $S_{X}$, typically given in the form
$\dot{\Sigma}\geq\frac{2J_{X}^{2}}{S_{X}}.$ (1)
Incorporating the information of fluctuations explicitly, it provides a universal finite lower bound that tightens the second law, $\dot{\Sigma}\geq 0$.
While extending the TUR to the quantum regime is a crucial problem gathering much attention recently [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], there is a fundamental issue of how to evaluate dynamical fluctuations. TURs usually involve fluctuations in fluxes or observables’ change [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], hence they require statistics more than single time points. However, in quantum mechanics, physical quantities at different times do not commute in the Heisenberg sense, which makes fluctuations elusive.
Conventionally, there are two approaches trying to characterize quantum fluctuations: one is based on the full counting statistics or continuous measurement [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which identifies jumps as exchanging charges with the environment or detections of a signal in an experiment [37]. However, it cannot explore fluctuations of more general observables that are not directly related to the jumps. The other approach utilizes the two-point measurement [14, 15, 16, 17], which, however, ends up with discarding coherence regarding the measured observable we are interested in, because it includes invasive measurement steps [38]. Hence, these methods can fail to capture all the quantum effects on thermodynamic trade-off relations.
Our idea to overcome this difficulty is to focus on quasiprobabilities to fully describe the quantum statistics of observables that are not detected by jumps and may exhibit non-commutativity. Quasiprobabilities are quantum extensions of the classical joint probabilities, and their peculiar behaviors have been attracting significant attention in recent years [39, 40, 41]. While the classical probabilities satisfy positivity and linearity and lead to correct marginals, any quantum counterparts are known to violate at least one of these three properties [42, 40]. As the violation of positivity is connected to genuine quantumness called contextuality [43, 44], quasiprobabilities that may take negative values have recently been studied extensively [45, 46, 47, 48, 49, 50, 51].
Figure 1: (a) In open systems, we cannot reduce two costs simultaneously: irreversibility, quantified by the entropy production rate $\dot{\Sigma}$, and fluctuations. That is represented by a universal trade-off relation called the thermodynamic uncertainty relation (TUR), where the product between $\dot{\Sigma}$ and the dynamical fluctuation $S_{X}$ of a physical quantity $X$ is bounded by a current strength $J_{X}$. (b) We prove a quantum TUR with $S_{X}=m_{X}$ [Eq. (10)], where the observable’s dynamical fluctuation is quantified by the quasiprobability, a quantum extension of the classical joint probability that may take negative values.
In this Letter, we derive a quantum TUR by quantifying the dynamical fluctuation of an observable with such a quasiprobability for the first time (see Fig. 1). We consider Markovian open quantum systems and show that the short-time variance of an observable’s change, as assessed by the Terletsky–Margenau–Hill quasiprobability [52, 53], provides a short-time TUR bound on the EPR for any state and observable. Moreover, on the basis of the non-classicality of quasiprobability, we elucidate criteria required for the anomalously large fluctuation absent in classical systems. Combined with our TUR, the criteria serve as a more fundamental, basis-independent condition for the dissipationless current [54, 55] than the abundance of the quantum coherence indicated in Ref. [54].
Setup.— We consider completely positive Markovian quantum dynamics, which are generally described by the quantum master equation [56, 57, 58]
$\frac{d\rho}{dt}=\mathcal{L}(\rho)\coloneqq-i[H,\rho]+\mathcal{D}(\rho),$ (2)
$\mathcal{D}(\rho)\coloneqq\sum_{k}L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\},$ (3)
where $\rho$ is the density operator, $H$ the Hamiltonian, $L_{k}$ jump operators, $\cdot$$\cdot$] the commutator, and $\{\cdot,\cdot\}$ the anticommutator. We refer to $\mathcal{L}$ and $\mathcal{D}$ as the Liouvillian and the dissipator, respectively. Given the initial data $\rho_{0}$, Eq. (2) is formally solved by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$. We indicate the adjoint of $\mathcal{L}$ and $\mathcal{D}$ by $\mathcal{L}^{\dagger}$ and $\mathcal{D}^{\dagger}$. For an observable (Hermitian operator) $X$, we describe the eigendecomposition as $X=\sum_{x}x\Pi_{x}$ with eigenvalues $x$ and projectors $\Pi_{x}$. Its expectation value at time $t$ is given by $\langle X\rangle_{t}=\mathop{\mathrm{tr}}\nolimits(X\rho(t))=\sum_{x}xp(x,t)$, where $p(x,t)=\mathop{\mathrm{tr}}\nolimits(\Pi_{x}\rho(t))$. We divide the time derivative into the Hamiltonian part $J_{X}^{\mathrm{H}}(t)=i\langle[H,X]\rangle_{t}$ and the dissipative part $J_{X}^{\mathrm{d}}(t)=\langle\mathcal{D}^{\dagger}(X)\rangle_{t}$ as $d_{t}\langle X\rangle_{t}=J_{X}^{\mathrm{H}}(t)+J_{X}^{\mathrm{d}}(t)$.
To discuss multi-time statistics of $X$ in the quantum dynamics beyond the single-time one $p(x,t)$, we introduce the quasiprobability [39, 40, 41]. In particular, we focus on the two-time statistics given by the Terletsky–Margenau–Hill (TMH) quasiprobability [52, 53], which is defined by
$q(y,t+\Delta t;x,t)\coloneq\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\Big{(}\big{\{}e^{\mathcal{L}^{\dagger}\Delta t}\Pi_{y},\Pi_{x}\big{\}}\rho(t)\Big{)}.$ (4)
The TMH quasiprobability is the real part of the Kirkwood–Dirac quasiprobability [59, 60]. While its marginals provide the single-time distributions as $\sum_{y}q(y,t+\Delta t;x,t)=p(x,t)$ and $\sum_{x}q(y,t+\Delta t;x,t)=p(y,t+\Delta t)$, it can become negative, unlike the classical joint probability distribution, which is an emergence of genuine quantumness [43, 44].
Analogously to the classical case [61], we define the moments of “change in XX” by
$\big{\langle}(\Delta X)^{n}\big{\rangle}_{t,t+\Delta t}\coloneqq\sum_{x,y}(x-y)^{n}q(y,t+\Delta t;x,t).$ (5)
It can be computed from the moment generating function $\mathsf{G}_{t}(\lambda,\Delta t)$ as $\langle(\Delta X)^{n}\rangle_{t,t+\Delta t}=(-i\partial_{\lambda})^{n}\mathsf{G}_{t}(\lambda,\Delta t)|_{\lambda=0}$, where $\mathsf{G}_{t}(\lambda,\Delta t)$ is defined as
$\mathsf{G}_{t}(\lambda,\Delta t)\coloneqq\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\Big{(}\big{\{}e^{\mathcal{L}^{\dagger}\Delta t}e^{i\lambda X},e^{-i\lambda X}\big{\}}\rho(t)\Big{)}.$ (6)
Theoretical studies have revealed that the moment generating function can be directly measured by devising an interferometric scheme [40, 39], which has been experimentally realized for $X=H$ [62]. Thus, $\langle(\Delta X)^{n}\rangle_{t,t+\Delta t}$ is not only theoretically but also practically a reasonable measure of the dynamical fluctuation in a quantum object $X$. Additionally, as shown in End Matter, $\mathsf{G}_{t}(\lambda,\Delta t)$ can be obtained by appropriately “quantizing” the classical counterpart.
We further define the short-time moments $m_{X}^{(n)}(t)$ by $m_{X}^{(n)}(t)\coloneqq\lim_{\Delta t\to 0}\langle(\Delta X)^{n}\rangle_{t,t+\Delta t}/\Delta t$. We write $m_{X}^{(2)}(t)$ simply as $m_{X}(t)$ and call it the short-time fluctuation of $X$ because the case $n=2$ will be of the most importance due to its direct connection to the variance, $m_{X}^{(2)}(t)=\lim_{\Delta t\to 0}\mathrm{Var}(\Delta X)/\Delta t$, where $\mathrm{Var}(\Delta X)\coloneqq\langle(\Delta X)^{2}\rangle_{t,t+\Delta t}-(\langle\Delta X\rangle_{t,t+\Delta t})^{2}$. The limit is always well defined because the TMH quasiprobability is expanded as $q(y,t+\Delta t;x,t)=\delta_{xy}p(x,t)+T_{yx}\big{(}\rho(t)\big{)}\Delta t+O(\Delta t^{2})$ with
$T_{yx}(\rho)\coloneqq\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\big{(}\{\mathcal{L}^{\dagger}\Pi_{y},\Pi_{x}\big{\}}\rho\big{)}$ (7)
when $\Delta t$ goes to zero. Because $\mathcal{L}^{\dagger}(I)=0$, we have $\sum_{y}T_{yx}(\rho)=0$. The short-time moments are given by $T_{yx}$ as
$m_{X}^{(n)}(t)=\sum_{x,y}(x-y)^{n}T_{yx}(\rho(t)).$ (8)
We refer to $T_{yx}$ as the flux from $x$ to $y$ in analogy to the classical Markov processes. Indeed, if $\rho(t)$ commutes with $X$ and $X$ has no degeneracy, then $\rho$ is diagonalized as $\rho(t)=\sum_{x}p(x,t)\Pi_{x}$ with $\Pi_{x}=|x\rangle\langle x|$ and we have $T_{yx}(\rho(t))=R_{yx}p(x,t)$ with $R_{yx}=\sum_{k}|\langle y|L_{k}|x\rangle|^{2}$. Now, $p(x,t)$ and $R_{yx}$ can be interpreted respectively as the occupation probability of the classical state labeled by $x$ and the transition rate from state $x$ to state $y$ [61, 63]. The diagonal elements $T_{xx}(\rho(t))$ also read $R_{xx}p(x,t)$, where $R_{xx}=-\sum_{x^{\prime}(\neq x)}\sum_{k}|\langle x^{\prime}|L_{k}|x\rangle|^{2}$, and $R_{xx}$ corresponds to the escape rate from state $x$. Now that, even in the general (quantum) cases, we call the sum $\bar{\lambda}(\rho)\coloneqq-\sum_{x}T_{xx}(\rho)$ the average escape rate.
Finally, we make a few assumptions to discuss thermodynamics. We assume that each jump $k$ has a unique counterpart k-k and that jump $k$ induces an entropy current $s_{k}$ such that $s_{-k}=-s_{k}$. The jump operators are postulated to satisfy the local detailed balance $L_{k}=e^{s_{k}/(2k_{\mathrm{B}})}L_{-k}^{\dagger}$ [64, 65, 66]. Then, we can define the entropy production rate by
$\dot{\Sigma}\big{(}\rho(t)\big{)}\coloneq k_{\mathrm{B}}\frac{d}{dt}S(t)+\sum_{k}s_{k}\mathop{\mathrm{tr}}\nolimits\big{(}L_{k}^{\dagger}L_{k}\rho(t)\big{)},$ (9)
where $S(t)\coloneq-\mathop{\mathrm{tr}}\nolimits(\rho(t)\ln\rho(t))$ is the von Neumann entropy. Hereafter, we set the Boltzmann constant $k_{\mathrm{B}}$ to one for simplicity.
Main result.— The following inequality is our main result:
$\dot{\Sigma}\big{(}\rho(t)\big{)}\geq\frac{2|J_{X}^{\mathrm{d}}(t)|^{2}}{m_{X}(t)}$ (10)
for any state $\rho(t)$ and observable $X$. The numerator is the dissipative part of the time derivative of $\langle X\rangle_{t}$, so it quantifies the changing rate of $X$ due to the dissipative dynamics. On the other hand, the denominator, the short-time fluctuation, represents the dynamical fluctuation of $X$ as evaluated by the TMH quasiprobability. Thus, the inequality represents the universal trade-off between dissipation and fluctuations, i.e., the TUR [2, 3]. Specifically, it generalizes the so-called short-time TURs in classical systems [67, 68], which have been utilized in estimating the EPR in classical systems [69, 70, 68, 71, 72]. This connection will be clear if we assume that $\rho$ commutes with $X$; then, the numerator and denominator read $J_{X}^{\mathrm{d}}(t)=d_{t}\langle X\rangle_{t}$ and $m_{X}(t)=\sum_{x,y}(x-y)^{2}R_{yx}p(x,t)$, which appear in the short-time TUR presented in Ref. [68]. The TUR indicates that increasing the short-time fluctuation can lead to the reduction of the lower bound of dissipation. As will be discussed in detail later, the quasiprobabilistic perspective has the practical advantage of clarifying physical requirements for such an increase in $m_{X}$.
Let us prove Eq. (10). We employ the following inequality derived by one of the present authors in Ref. [73]:
$\dot{\Sigma}\big{(}\rho(t)\big{)}\geq\frac{|\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho(t)))|^{2}}{\mathfrak{D}_{X}(\rho(t))},$ (11)
where $\mathfrak{D}_{X}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits(\rho(\mathcal{D}^{\dagger}(X^{2})-\{\mathcal{D}^{\dagger}(X),X\}))$ is called the quantum diffusivity (the derivation is reviewed in Supplemental Material [74]). It is easy to see $\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho(t)))=J_{X}^{\mathrm{d}}(t)$. The nontrivial point is that we can also associate the denominator $\mathfrak{D}_{X}(\rho(t))$ with $m_{X}(t)$. First, by expanding the definition, we find
$m_{X}(t)$ $=\frac{1}{2}\sum_{x,y}(y-x)^{2}\mathop{\mathrm{tr}}\nolimits(\{\mathcal{L}^{\dagger}\Pi_{y},\Pi_{x}\}\rho(t))$ (12)
$=\mathop{\mathrm{tr}}\nolimits\big{(}\mathcal{L}^{\dagger}(X^{2})\rho(t)\big{)}-\mathop{\mathrm{tr}}\nolimits\big{(}\{\mathcal{L}^{\dagger}(X),X\}\rho(t)\big{)}.$ (13)
Moreover, by using an identity $\{[H,X],X\}=[H,X^{2}]$, we can remove the Hamiltonian part from $\mathcal{L}^{\dagger}$. Therefore, we find the equality
$\mathfrak{D}_{X}\big{(}\rho(t)\big{)}=\frac{1}{2}m_{X}(t).$ (14)
Combining it with Eq. (11), we obtain the TUR (10).
Comparison with existing TURs.—We compare our TUR (10) with existing quantum TURs. First, we stress the difference from widely studied methods, the full counting statistics (FCS) approach [37] and the two-point measurement (TPM) approach [38]. In short, our result is complementary to those previous ones. While the TURs based on the FCS [6, 7, 8, 9, 10, 11, 12, 13] deal with current observables associated with the counting of jumps, our TUR is given by the fluctuations of intrinsic observables, which do not have to be connected to jumps. It is noteworthy that in classical Markov jump processes, an intrinsic observable’s change is entirely given by keeping track of jumps, because every physical quantity takes determined values in classical states [61]. However, quantum states linked by quantum jumps are not eigenstates of every intrinsic observable (they may not be even for the Hamiltonian; see Model A in [37]). In End Matter, we discuss a nontrivial intersection between the FCS and quasiprobabilities when $X,L_{k}]=w_{k}L_{k}$ holds for an observable $X$ with certain weights $w_{k}$.
The TPM method, adopted to quantify dynamical fluctuations in Refs. [14, 15, 16, 17], inevitably leads to decoherence in the initial density operator due to invasive initial measurement, despite having a clear experimental perspective. On the other hand, in our TUR, every initial coherence is incorporated owing to the employment of the quasiprobability, while maintaining experimental access by the interferometric method [40, 39, 62].
Finally, we mention the connection to the “TUR” [Eq. (11)] given in Ref. [73]. Despite the apparent similarity, our TUR has a substantial advantage over the previous result. As is clear from the proof, our TUR is derived by combining their result and the equality (14). The latter equality gives a clear statistical meaning to $\mathfrak{D}_{X}(\rho)$, which was interpreted as a fluctuation measure only in classical cases in Ref. [73]. Moreover, as discussed below, the quasiprobabilistic perspective brought by Eq. (14) enables us to understand a non-classical suppression of dissipation through non-classical behaviors of quasiprobabilities, which cannot be understood from Eq. (11) alone.
Anomalous scaling via non-classicality of quasiprobability.— We next explore the ramifications of non-classicality on the thermodynamic constraint. As will be shown, the short-time fluctuation $m_{X}(t)$ can exhibit an anomalous scaling when $X$ is highly degenerate, which is forbidden in classical systems. This scaling makes the TUR bound small and allows for reducing entropy production beyond the classical limit. While the reduction has recently been attributed to quantum coherence in $\rho$ [54], we discuss that non-classical behaviors of the TMH quasiprobability are more fundamental.
First, we review the discussion in Ref. [54]. The authors derived an inequality similar to Eq. (10) for $X=H$, the Hamiltonian, where they used the coherence in $\rho$ regarding a specific eigenbasis of $H$ instead of the short-time fluctuation $m_{H}$. They showed that when $H$ is highly degenerate, dissipationless heat current can be realized, as $\dot{\Sigma}(\rho(t))=O(1)$ and $|d_{t}\langle H\rangle_{t}|=|J_{H}^{\mathrm{d}}(t)|=O(N)$, where $N$ is the number of degeneracy. In our terminology, this dissipationless current is enabled by $m_{H}(t)$ scaling at $O(N^{2})$, while they attributed it to the coherence’s $O(N)$ scaling.
To clarify what is crucial for the dissipationless current, we formulate the problem in a general way. We assume that the eigenvalues of $X$ are $N$-fold degenerate and examine conditions on $\rho(t)$ for $m_{X}(t)$ to scale more rapidly than the classical limitation $m_{X}(t)=O(N)$ (detailed below). We specifically refer to the growth of the short-time fluctuation faster than $O(N)$ as anomalous scaling. When an eigenbasis for eigenvalue $x_{s}$ is given as $\{|s,j\rangle\}_{j=1}^{N}$, $X$ is expanded as $X=\sum_{s,j}x_{s}\Pi_{s,j}$ with $\Pi_{s,j}=|s,j\rangle\langle s,j|$. We can define the flux from $(s,j)$ to $(s^{\prime},j^{\prime})$ as $T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits(\{\mathcal{L}^{\dagger}\Pi_{s^{\prime},j^{\prime}},\Pi_{s,j}\}\rho)$ and the integrated fluxes as
$\mathcal{T}_{s^{\prime}s}(\rho)$ $\coloneqq\begin{cases}\sum_{j,j^{\prime}}T_{s^{\prime}j^{\prime}sj}(\rho)&(s\neq s^{\prime}),\\ \sum_{j,j^{\prime}(\neq j)}T_{sj^{\prime}sj}(\rho)&(s=s^{\prime}).\end{cases}$ (15)
While $T_{s^{\prime}j^{\prime}sj}$ and $\mathcal{T}_{ss}$ are basis-dependent, $\mathcal{T}_{s^{\prime}s}$ are not if $s\neq s^{\prime}$. We further define the average escape rate $\bar{\lambda}(\rho)=-\sum_{s,j}T_{sjsj}(\rho)$. They satisfy the equality $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)-\bar{\lambda}(\rho)=0$. Because the time dependence is not relevant hereafter, we write $m_{X}(\rho)$ to indicate $m_{X}(t)$ when $\rho(t)=\rho$. The short-time fluctuation is given solely by the integrated fluxes as $m_{X}(\rho)=\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}\mathcal{T}_{s^{\prime}s}(\rho)$.
We first demonstrate $m_{X}(\rho)=O(N)$ in the classical cases. To this end, we define an eigenbasis to be classical if it satisfies that for each $(s,j,s^{\prime},j^{\prime})$, the number of jumps such that $|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|\neq 0$ is $O(1)$, and that $|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|=O(1)$. For a classical eigenbasis, we find $R_{s^{\prime}j^{\prime}sj}=\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(1)$ and $R_{sjsj}=\sum_{s^{\prime},j^{\prime}(\neq s,j)}\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(N)$. Therefore, if $\rho$ has no coherence with respect to the classical eigenbasis (i.e., $\rho=\sum_{s,j}p_{sj}\Pi_{s,j}$ holds), we obtain the scaling $\bar{\lambda}(\rho)=-\sum_{s,j}R_{sjsj}p_{sj}=O(N)$ because $p_{sj}\geq 0$ and $\sum_{s,j}p_{sj}=1$. This scaling is intuitive as the target of escape grows by at most $O(N)$. Because the absence of coherence in $\rho$ also leads to the positivity of $T_{s^{\prime}j^{\prime}sj}(\rho)$, we can conclude $m_{X}(\rho)=O(N)$ from
$m_{X}(\rho)=\sum_{s,s^{\prime}}(x_{s}-x_{s}^{\prime})^{2}\mathcal{T}_{s^{\prime}s}(\rho)$
$\leq\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$
$=\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}\bar{\lambda}(\rho)=O(N).$
Here, we have assumed that $\mathcal{T}_{s^{\prime}s}(\rho)\neq 0$ implies $|x_{s}-x_{s^{\prime}}|=O(1)$, which is true if, for example, there is a degeneracy-independent threshold $\Lambda$ such that $|x_{s}-x_{s^{\prime}}|>\Lambda$ implies $\langle s^{\prime},j^{\prime}|H|s,j\rangle=0$ and $\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle=0$ (namely, transitions leading to a drastic change in $X$ are forbidden).
The above discussion shows that if $\rho$ has no coherence regarding the classical eigenbases, $m_{X}(\rho)$ grows by at most $O(N)$. On the other hand, if we adopt a non-classical eigenbasis, the coherence tells us nothing. For example, consider $\rho$ expanded as $\rho=\sum_{s}p_{s}|s,*\rangle\langle s,*|$ with $|s,*\rangle$ such that $X|s,*\rangle=x_{s}|s,*\rangle$ and $|\langle s^{\prime},*^{\prime}|L_{k}|s,*\rangle|=O(N^{\alpha})$ with $\alpha>1/2$. Then, $\bar{\lambda}$ grows by $O(N^{2\alpha})>O(N)$ and an anomalous scaling can occur even though $\rho$ has no coherence. We will exemplify this drawback of coherence-based consideration through an example later.
The next statement is our second main result, which provides another viewpoint to examine the anomalous scaling: if neither of the following conditions is satisfied for some eigenbasis, $m_{X}(\rho)$ does not exhibit anomalous scaling:
(Q1) $\exists s,s^{\prime},\;\;\lim_{N\to\infty}\frac{\mathcal{T}_{s^{\prime}s}(\rho)}{N}=-\infty,$
(Q2) $\lim_{N\to\infty}\frac{\bar{\lambda}(\rho)}{N}=\infty.$
The first condition (Q1) represents significant negativity in the fluxes, which results in negative quasiprobability because then $T_{s^{\prime}j^{\prime}sj}(\rho)<0$ for several $j,j^{\prime}$ and
$q((s^{\prime},j^{\prime}),t+\Delta t;(s,j),t)$
$=T_{s^{\prime}j^{\prime}sj}(\rho(t))\Delta t+O(\Delta t^{2})<0.$ (16)
Note that the negativity of the quasiprobability also indicates the genuine quantumness called contextuality [43, 44]. The other one (Q2) is also non-classical because it means the chance of escape grows more rapidly than the number of evacuation targets increases. The statement shows that either of these quantumness conditions has to hold in any basis for the anomalous scaling to occur. Such a property stands in contrast to the existence of coherence, which does not need to hold in every basis. We depict their relationship with anomalous scaling in Fig. 2.
Figure 2: Hierarchy of conditions. Condition (Q1) or (Q2) is always necessary for $m_{X}(\rho)$ to scale anomalously, as is prohibited in classical systems. (a) For classical eigenbases, these conditions are sufficient for the state to have coherence. Thus, this time, coherence is necessary for the anomalous scaling. Density matrix $\rho^{+}$, discussed in Example, has a large coherence regarding a classical basis to exhibit anomalous scaling. (b) If we take a non-classical eigenbasis, coherence regarding that basis loses its connection to the scaling ($\rho^{+}$ may become “incoherent”). On the other hand, our two conditions remain relevant. In both cases, there can be a state $\rho^{-}$ that has as much coherence as $\rho^{+}$ but does not satisfy (Q1) or (Q2) and thus does not show anomalous scaling.
Let us prove the claim. First, because $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\bar{\lambda}(\rho)$, the negation of (Q2) implies that $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$ is $O(N)$. Then, split the summation as $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\mathcal{T}^{+}-\mathcal{T}^{-}$ with $\mathcal{T}^{+}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\geq 0}\mathcal{T}_{s^{\prime}s}(\rho)$ and $\mathcal{T}^{-}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)<0}|\mathcal{T}_{s^{\prime}s}(\rho)|$. Due to the negation of (Q1), the negative part $\mathcal{T}^{-}$ is at most $O(N)$. Therefore, the positive part $\mathcal{T}^{+}$ can also be at most $O(N)$. Finally, $m_{X}(\rho)$ is bounded as
$m_{X}(\rho)$ $\leq\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}|\mathcal{T}_{s^{\prime}s}(\rho)|$
$\leq\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}(\mathcal{T}^{+}+\mathcal{T}^{-}),$
and thus is of order $N$.
Example.— We illustrate the above discussion through the model used in Ref. [54]. Its Hamiltonian $H$ has two $N$-fold degenerate eigenvalues, 0 and $\omega$. The jump operators represent simultaneous jumps between energy levels; with an eigenbasis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$ such that $H|g,j\rangle=0$ and $H|e,j\rangle=\omega|e,j\rangle$, they are defined as $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. Anomalous scaling $m_{H}(\rho)=O(N^{2})$ is observed if we set $\rho=\rho^{+}=p_{g}|g,+\rangle\langle g,+|+p_{e}|e,+\rangle\langle e,+|$ with $|s,+\rangle=\frac{1}{\sqrt{N}}\sum_{j}|s,j\rangle$ for $s=e,g$ (see Supplemental Material [74]). The state $\rho^{+}$ has a large $l_{1}$-coherence $C_{l_{1}}(\rho^{+})=O(N)$, defined as $C_{l_{1}}(\rho)=\sum_{s,j,s^{\prime},j^{\prime}|(s,j)\neq(s^{\prime},j^{\prime})}|\langle s,j|\rho|s^{\prime},j^{\prime}\rangle|$, to which Ref. [54] attributed the dissipationless current. Now, note that the basis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$, defining the $l_{1}$-coherence, is classical since there are only two jumps regardless of $N$ and, for example, $\langle e,j^{\prime}|L_{+}|g,j\rangle=\sqrt{\gamma_{+}}=O(1)$. On the other hand, if we consider a non-classical basis including $|g,+\rangle$ and $|e,+\rangle$ (note that $\langle e,+|L_{+}|g,+\rangle=\sqrt{\gamma_{+}}N$), the state $\rho^{+}$ has no coherence regarding this basis. This fact shows that $m_{H}(\rho^{+})$ may exhibit anomalous scaling (that is, we may have dissipationless current) even if $\rho^{+}$ has no coherence if the referred basis is not classical. Nonetheless, we see that $\rho^{+}$ fulfills either (Q1) or (Q2) for any eigenbasis, from the general discussion.
In addition, we can construct a state $\rho^{-}$ that has as much coherence regarding the classical eigenbasis as $\rho^{+}$ but does not yield dissipationless current, violating both of the quantumness conditions, (Q1) and (Q2). Such a state is constructed by replacing $|s,+\rangle$ in the definition of $\rho^{+}$ with $|s,-\rangle=\frac{1}{\sqrt{N}}\sum_{j}(-1)^{j}|s,j\rangle$. While it has the same amount of $l_{1}$-coherence as $\rho^{+}$, the integrated fluxes and average escape rate read $\mathcal{T}_{s^{\prime}s}(\rho^{-})=\gamma_{s}p_{s}\chi(N)$ for $s\neq s^{\prime}$, $\mathcal{T}_{ss}(\rho^{-})=\frac{\gamma_{s}p_{s}}{2}(N-\chi(N))$, and $\bar{\lambda}(\rho^{-})=\frac{\gamma_{+}p_{g}+\gamma_{-}p_{e}}{2}(N+\chi(N))$, where $\gamma_{g}=\gamma_{+}$, $\gamma_{e}=\gamma_{-}$, and $\chi(N)$ is the remainder when $N$ is divided by $2$ [74]. Therefore, $\rho^{-}$ does not satisfy either (Q1) and (Q2), and leads to $m_{H}(\rho^{-})=O(1)$. Hence, anomalous dissipationless current does not occur for $\rho^{-}$, which cannot be understood by looking at the coherence alone.
Discussion.— We have explored fundamental connections between quantum thermodynamics and quasiprobability; in particular, we have derived a quantum TUR evaluating fluctuations through the TMH quasiprobability. The TUR is valid for any state $\rho$ and observable $X$, as long as the dynamics satisfy a few assumptions necessary to discuss thermodynamics. As discussed, our TUR is complementary to existing TURs, which adopt different approaches when quantifying fluctuations. We have also demonstrated that when discussing anomalous scaling of fluctuations in the TUR, which brings reduction of dissipation beyond classical limitations, non-classical behaviors of quasiprobability are more essential than quantum coherence.
To the best of our knowledge, our study is the first to bridge quasiprobabilities and the thermodynamic trade-off relations, recently discovered in stochastic thermodynamics [75]. In addition to the TUR, several fundamental relations have been unveiled in the field, such as thermodynamic speed limits [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86], bounds on asymmetry [87, 88, 89, 90], and fluctuation-response inequalities [91, 92, 93]. We expect that quasiprobabilities will help explore non-classical features in their quantum extensions.
Acknowledgements.
Acknowledgments.— K.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. K.Y. and R.H. are supported by JST ERATO Grant No. JPMJER2302, Japan.
Supplemental Material
.3 Derivation of Eq. (11)
We concisely review the derivation of Eq. (11) in Ref. [73]. For the proof of unproven facts, please see that paper.
The fundamental strategy is as follows: 1) rewrite the entropy production rate as an inner product, 2) rewrite it as a squared norm, and 3) apply the Cauchy–Schwarz inequality. First, notice that for each pair of jumps (k,k)(k,-k), we can take coefficients $\gamma_{\pm k}>0$ and operators $\tilde{L}_{\pm k}$ such that $L_{\pm k}=\sqrt{\gamma_{\pm k}}\tilde{L}_{\pm k}$ and $\tilde{L}_{-k}=\tilde{L}_{k}^{\dagger}$. The local detailed balance imposes $\gamma_{k}/\gamma_{-k}=e^{s_{k}}$ ($k_{\mathrm{B}}$ is set to be unity). Then, we define
$\mathbb{J}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&J_{-k}(\rho)\\ J_{k}(\rho)&O\end{pmatrix}$, (35)
$\mathbb{F}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&F_{-k}(\rho)\\ F_{k}(\rho)&O\end{pmatrix}$, (36)
where
$J_{k}(\rho)$ $\coloneqq\frac{1}{2}(\gamma_{k}\tilde{L}_{k}\rho-\gamma_{-k}\rho\tilde{L}_{k}),$ (37)
$F_{k}(\rho)$ $\coloneqq s_{k}\tilde{L}_{k}+[\tilde{L}_{k},\ln\rho].$ (38)
Now, $\mathbb{J}_{k}(\rho)$ and $\mathbb{F}_{k}(\rho)$ are anti-Hermitian operators on $\mathbb{C}^{2}\otimes\mathcal{H}$, where $\mathcal{H}$ is the original Hilbert space. By aligning them, we further define
$\mathbb{J}(\rho)\coloneqq\bigoplus_{k}\mathbb{J}_{k}(\rho),\quad\mathbb{F}(\rho)\coloneqq\bigoplus_{k}\mathbb{F}_{k}(\rho),$ (39)
where the direct sum is taken over the pairs of jumps (thus, if $k$ is counted, k-k is not). They become operators on $\mathfrak{h}\otimes\mathcal{H}$, where $\mathfrak{h}$ is a complex vector space whose dimension is equal to the number of jumps. Then, the entropy production rate (EPR) defined in Eq. (9) is given by
$\dot{\Sigma}(\rho)=\langle\mathbb{J}(\rho),\mathbb{F}(\rho)\rangle,$ (40)
where $\langle\cdot,\cdot\rangle$ is the Hilbert–Schmidt inner product, defined as $\langle A,B\rangle=\mathop{\mathrm{tr}}\nolimits(A^{\dagger}B)$.
Next, we introduce a super-operator $\nabla_{\mathbb{L}}$ that maps operators on $\mathcal{H}$ to those on $\mathfrak{h}\otimes\mathcal{H}$. It is defined by
$\nabla_{\mathbb{L}}A\coloneqq[I_{\mathfrak{h}}\otimes A,\mathbb{L}],$ (41)
where $I_{\mathfrak{h}}$ is the identity operator on $\mathfrak{h}$ and
$\mathbb{L}\coloneqq\bigoplus_{k}\begin{pmatrix}O&\tilde{L}_{-k}\\ \tilde{L}_{k}&O\end{pmatrix}$. (42)
It acts as a gradient, and its adjoint turns $\mathbb{J}(\rho)$ into the dissipator as $\nabla_{\mathbb{L}}^{\dagger}\mathbb{J}(\rho)=\mathcal{D}(\rho)$.
Finally, for a positive operator $G$, we introduce a super-operator
$\mathcal{M}_{G}(A)\coloneqq\int_{0}^{1}G^{s}XG^{1-s}ds,$ (43)
which satisfies $\mathcal{M}_{\Gamma\otimes\rho}(\mathbb{F}(\rho))=\mathbb{J}(\rho)$ with
$\Gamma=\bigoplus_{k}\begin{pmatrix}\gamma_{k}/2&0\\ 0&\gamma_{-k}/2\end{pmatrix}$. (44)
We can define an inner product $\langle A,B\rangle_{G}=\langle A,\mathcal{M}_{G}(B)\rangle$ and the induced norm $\lVert A\rVert_{G}=\sqrt{\langle A,A\rangle_{G}}$, which allows us to rewrite the EPR as
$\dot{\Sigma}(\rho)=\lVert\mathbb{F}(\rho)\rVert_{\Gamma\otimes\rho}^{2}.$ (45)
By applying the Cauchy–Schwarz inequality, we get
$\dot{\Sigma}(\rho)\lVert\nabla_{\mathbb{L}}X\rVert_{\Gamma\otimes\rho}^{2}\geq|\langle\mathbb{F}(\rho),\nabla_{\mathbb{L}}X\rangle_{\Gamma\otimes\rho}|^{2}.$ (46)
Because $\mathcal{M}_{\Gamma\otimes\rho}(\mathbb{F}(\rho))=\mathbb{J}(\rho)$ holds, the right-hand side leads to $\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))$. On the other hand, we can generally show $\langle A,\mathcal{M}_{G}(A)\rangle\leq\frac{1}{2}\mathop{\mathrm{tr}}\nolimits(A^{\dagger}\{G,A\})$; by a direct calculation, we can derive that this upper bound leads to $\mathfrak{D}_{X}(\rho)$ when $A=\nabla_{\mathbb{L}}X$ and $G=\Gamma\otimes\rho$ and obtain Eq. (11).
Appendix A Computational details of Example
We provide the details of the computation in the example in the main text. We aim to compute $\mathcal{T}_{s^{\prime}s}(\rho)$ and $\bar{\lambda}(\rho)$ for $\rho=\rho^{\pm}$ and show $m_{H}(\rho^{+})=O(N^{2})$ and $m_{H}(\rho^{-})=O(1)$. Here, the model we consider consists of Hamiltonian $H=\omega\sum_{j}|e,j\rangle\langle e,j|$ and jump operators $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. The states of interest $\rho^{\pm}$ are generated as $\rho^{\pm}=\sum_{s=g,e}p_{s}|s,\pm\rangle\langle s,\pm|$, where $|s,\pm\rangle=\frac{1}{\sqrt{N}}\sum_{j}(\pm 1)^{j}|s,j\rangle$ and $p_{s}$ satisfies $p_{s}\geq 0$ and $p_{g}+p_{e}=1$.
Because now we consider $X=H$, we need to compute
$T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\big{(}\{\mathcal{D}^{\dagger}\Pi_{s^{\prime}j^{\prime}},\Pi_{sj}\}\rho\big{)}.$ (47)
For convenience, we define
$\sigma_{s,j,j^{\prime}}$ $\coloneqq\frac{1}{2}\big{(}|s,j\rangle\langle s,j^{\prime}|+|s,j^{\prime}\rangle\langle s,j|\big{)},$ (48)
$\Lambda_{s,j}$ $\coloneqq\sum_{j^{\prime}}\sigma_{s,j,j^{\prime}}.$ (49)
They satisfy
$\sigma_{s,j,j^{\prime}}=\sigma_{s,j^{\prime},j},\;\;\sigma_{s,j,j}=\Pi_{s,j},$ (50)
$\mathop{\mathrm{tr}}\nolimits(\sigma_{s,j,j^{\prime}}\rho)=\mathrm{Re}\langle s,j|\rho|s,j^{\prime}\rangle,$ (51)
$\sum_{j}\Lambda_{s,j}=\sum_{j,j^{\prime}}|s,j\rangle\langle s,j^{\prime}|,$ (52)
$\frac{1}{2}\sum_{j^{\prime}}\{\Lambda_{s^{\prime},j^{\prime}},\Pi_{s,j}\}=\delta_{s,s^{\prime}}\Lambda_{s,j}.$ (53)
By using them, we obtain
$L_{+}^{\dagger}\Pi_{s,j}L_{+}$ $=\gamma_{+}\delta_{s,e}\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (54)
$L_{-}^{\dagger}\Pi_{s,j}L_{-}$ $=\gamma_{-}\delta_{s,g}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}},$ (55)
and
$L_{+}^{\dagger}L_{+}$ $=\gamma_{+}N\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (56)
$L_{-}^{\dagger}L_{-}$ $=\gamma_{-}N\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}.$ (57)
Then, we find
$\mathcal{D}^{\dagger}\Pi_{s,j}$ $=\delta_{s,e}\bigg{(}\gamma_{+}\sum_{j^{\prime}}\Lambda_{g,j^{\prime}}-\gamma_{-}N\Lambda_{e,j}\bigg{)}$
$+\delta_{s,g}\bigg{(}\gamma_{-}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}-\gamma_{+}N\Lambda_{g,j}\bigg{)}.$ (58)
Consequently, we get
$\frac{1}{2}\{\mathcal{D}^{\dagger}\Pi_{s^{\prime},j^{\prime}},\Pi_{s,j}\}$
$=\delta_{s^{\prime},e}\delta_{s,g}\gamma_{+}\Lambda_{g,j}-\frac{\delta_{s^{\prime},e}\delta_{s,e}}{2}\gamma_{-}N(\delta_{j,j^{\prime}}\Lambda_{e,j}+\sigma_{e,j,j^{\prime}})$
$+\delta_{s^{\prime},g}\delta_{s,e}\gamma_{-}\Lambda_{e,j}-\frac{\delta_{s^{\prime},g}\delta_{s,g}}{2}\gamma_{+}N(\delta_{j,j^{\prime}}\Lambda_{g,j}+\sigma_{g,j,j^{\prime}}).$ (59)
For density matrix $\rho$, the fluxes are given as
$T_{ej^{\prime}gj}(\rho)$ $=\gamma_{+}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle$ (60)
$T_{gj^{\prime}ej}(\rho)$ $=\gamma_{-}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle$ (61)
$T_{gj^{\prime}gj}(\rho)$ $=-\frac{\gamma_{+}N}{2}\bigg{(}\mathrm{Re}\langle g,j|\rho|g,j^{\prime}\rangle$
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle\bigg{)}$ (62)
$T_{ej^{\prime}ej}(\rho)$ $=-\frac{\gamma_{-}N}{2}\bigg{(}\mathrm{Re}\langle e,j|\rho|e,j^{\prime}\rangle$
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle\bigg{)}.$ (63)
For $\rho=\rho^{+}$ and $\rho^{-}$, because $\langle s,j|\rho^{+}|s,j^{\prime\prime}\rangle=p_{s}/N$ and $\langle s,j|\rho^{-}|s,j^{\prime\prime}\rangle=(-1)^{j+j^{\prime\prime}}p_{s}/N$, we find
$T_{ej^{\prime}gj}(\rho^{+})=p_{g}\gamma_{+},\;\;T_{ej^{\prime}gj}(\rho^{-})=(-1)^{j+1}p_{g}\gamma_{+}\frac{\chi(N)}{N},$ (64)
where we used $\sum_{j}(-1)^{j}=-\chi(N)$. Thus,
$\mathcal{T}_{eg}(\rho^{+})=p_{g}\gamma_{+}N^{2},\quad\mathcal{T}_{eg}(\rho^{-})=p_{g}\gamma_{+}\chi(N).$ (65)
Note that $\chi(N)^{2}=\chi(N)$. We also see
$T_{gj^{\prime}gj}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}(1+\delta_{j,j^{\prime}}N)$ (66)
$T_{gj^{\prime}gj}(\rho^{-})$ $=-\frac{(-1)^{j}\gamma_{+}p_{g}}{2}((-1)^{j^{\prime}}-\delta_{j,j^{\prime}}\chi(N)).$ (67)
Therefore, we can get
$\mathcal{T}_{gg}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}N(N-1),$ (68)
$\mathcal{T}_{gg}(\rho^{-})$ $=\frac{\gamma_{+}p_{g}}{2}(N-\chi(N))$ (69)
because
$\sum_{j,j^{\prime}(\neq j)}(-1)^{j}(-1)^{j^{\prime}}$ $=\bigg{(}\sum_{j}(-1)^{j}\bigg{)}^{2}-\sum_{j}(-1)^{2j}$ (70)
$=\chi(N)-N.$ (71)
The other integrated fluxes are given in the same way. Besides, we can compute the average escape rates as
$\bar{\lambda}(\rho^{+})$ $=\frac{\gamma_{+}p_{g}+\gamma_{-}p_{e}}{2}N(N+1),$ (72)
$\bar{\lambda}(\rho^{-})$ $=\frac{\gamma_{+}p_{g}+\gamma_{-}p_{e}}{2}(N+\chi(N)).$ (73)
Finally, from Eq. (65), we find
$m_{H}(\rho^{+})$ $=\omega^{2}(\gamma_{+}p_{g}+\gamma_{-}p_{e})N^{2}=O(N^{2}),$ (74)
$m_{H}(\rho^{-})$ $=\omega^{2}(\gamma_{+}p_{g}+\gamma_{-}p_{e})\chi(N)=O(1).$ (75)